how to do you solve systems of equation (-1,3)
y=-x+2 y=x+4 step-by-step showing work

Answer by indra89811(24)

We can solve these equations by Elimination method.
First, adding these two equations. we get
2y=6
Dividing both sides by 2,
y=6/2=3
So, y=3.
Now, plugging in y=3 in equation (2), we get
3=x+4
Subtracting 4 from both sides,
3-4=x
-1=x
Thus, the solution is (-1,3)
---------------------Peace out, hun :) --------------------